what is most nearly the cost of concrete to repair a bridge wing wall. the wing wall measurement are 20 feet 6 inches long by 7 feet 9 inches wide by 10 inches thick delivered to the site concrete is $ 81.00 per cubic yard including tax and must be ordered by the full yard
Answer by mananth(16946)

20 feet 6 inches long = 6.83 yards
by 7 feet 9 inches wide = 2.58 yards
by 10 inches =0.28 yards
So volume of concrete = 6.83*2.58*0.28 cu. yards=4.93 = 5 cu.yard
concrete is $ 81.00 per cubic yard
5*81 = cost of concrete= $405